# footnotetext: MSC2020: 53A04, 53A70 (primary), 53A17, 70B15, 37K25, 37K10, 35Q53 (secondary)

### v2

@@ 5.4. Potential function of curvature
A proof of Theorem 5.14 is given in Appendix G. The potential function simultaneously solves the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation.
@@
(5.34) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$ $=\frac{L_{2}(v,y)}{y}\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right),$
where $Q_{1},Q_{2},L_{1},L_{2}\in\mathbb{R}$ are defined by
(5.35) $Q_{1}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)},\quad Q_{2}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}\left(0;\frac{2i}{y}\right)},$
(5.36) $L_{1}(v,y)=d_{1}\left(\frac{v}{y};\frac{2}{y}\right)-d_{4}\left(\frac{v}{y};\frac{2}{y}\right)+Q_{1}(v,y)+Q_{2}(v,y),$
(5.37) $L_{2}(v,y)=d_{1}\left(\frac{v}{y};\frac{2}{y}\right)-d_{4}\left(\frac{v}{y};\frac{2}{y}\right)-Q_{1}(v,y)-Q_{2}(v,y).$
A proof of Theorem 5.15 is given in Appendix H.
The following shows two equations are related by a parameter shift.
Proposition 5.16.
By the transformation of parameters $v\rightarrow v+y$ and $r\rightarrow\frac{1}{2}-r$, the signed curvature angle $\kappa$, the torsion angles $\lambda$, the potential functions $\Theta$, the coefficients of the semi-discrete mKdV equation $L_{1}/y$ and of the semi-discrete sine-Gordon equation $L_{2}/y$, the normal $\widetilde{N}$ and the binormal $\widetilde{B}$ are transformed as
(5.38) $\kappa_{n}\left(t\right)$ $\rightarrow(-1)^{n+1}\kappa_{n}\left(t-\frac{y}{2}\right),\quad\lambda\rightarrow\pi+\lambda,$
(5.39) $\Theta_{n}\left(t\right)$ $\rightarrow(-1)^{n}\Theta_{n}\left(t-\frac{y}{2}\right),$
(5.40) $\frac{L_{1}}{y}$ $\rightarrow\frac{L_{2}}{y},\quad\frac{L_{2}}{y}\rightarrow\frac{L_{1}}{y},$
(5.41) $\widetilde{N}_{n}(t)$ $\rightarrow(-1)^{n+1}\widetilde{N}_{n}\left(t-\frac{y}{2}\right),\quad\widetilde{B}_{n}(t)\rightarrow(-1)^{n+1}\widetilde{B}_{n}\left(t-\frac{y}{2}\right).$
A proof of Proposition 5.16 is given in Appendix I. The transformation $\Theta_{n}\rightarrow\widetilde{\Theta}_{n}:=(-1)^{n}\Theta_{n}$ exchanges a solution of the semi-discrete potential mKdV equation (5.33) with one of the semi-discrete sine-Gordon equation (5.34). Proposition 5.16 shows that this equation-level transformation is also achieved at the solution level by transforming the parameters.
@@
(5.42) $\frac{d}{dt}\Theta_{n}=4\beta^{2}\eta_{3}\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right),$
@@
which is also called the semi-discrete potential mKdV equation [14, 13, 6]. In other words, the potential function $\Theta$ defined by (5.31) is a simultaneous solution of three differential-difference equations, (5.33), (5.34) and (5.42). It does not usually happen that these equations are simultaneously solved. For example, it can be checked that the 2 and 3-soliton solutions of the semi-discrete sine-Gordon equation constructed in [6, 7] are not solutions of the semi-discrete potential mKdV equation.
The following describes an explicit relation between our construction and the previous work [19] reviewed in §3.2.
Proposition 5.18.
When $C_{+}$ and $\Gamma_{+}$ are chosen as in (5.18), we have
(5.43) $\frac{L_{1}(v,y)}{y}\frac{1}{1+\cos\lambda}=i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}=\rho_{+}.$
Similarly, when $C_{-}$ and $\Gamma_{-}$ are chosen as in (5.18), we have
(5.44) $\frac{L_{2}(v,y)}{y}\frac{1}{1-\cos\lambda}=-i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{2}}{\vartheta_{4}(r)^{2}\vartheta_{2}(r)^{2}}=\rho_{-}.$

### v3

@@ 5.4. Potential function of curvature
Proof.
From (A.17), (A.18) and (A.19), we see that
(5.151) $\begin{split}&\kappa_{n}=2\arctan\left(-i\frac{\vartheta_{1}(iv;iy)\vartheta_{4}\left(iv\left(n-\frac{1}{2}\right)+it;iy\right)}{\vartheta_{3}(iv;iy)\vartheta_{2}\left(iv\left(n-\frac{1}{2}\right)+it;iy\right)}\right)\\ &=2\arctan\left(-i\frac{\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}-\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}}{1-\frac{\vartheta_{1}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)\vartheta_{1}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}{\vartheta_{2}\left(i\frac{v}{2}\left(n+\frac{1}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)\vartheta_{2}\left(i\frac{v}{2}\left(n-\frac{3}{2}\right)+\frac{it}{2};\frac{iy}{2}\right)}}\right)\\ &=2\arctan\left(\frac{\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}-\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}}{1+\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}}\right)\\ &=2\arctan\left(\frac{\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}\right)\\ &-2\arctan\left(\frac{\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{3}{2}\right)+\frac{t}{y};\frac{2i}{y}\right)}\right)\\ &=\frac{\Theta_{n+1}-\Theta_{n-1}}{2},\end{split}$
holds. ∎
The potential function simultaneously solves the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation, establishing an explicit relation between our construction and the previous work [21] reviewed in §3.2. Since the semi-discrete sine-Gordon equation can be solved by almost the same calculation, we will solve only the semi-discrete potential mKdV equation here. We construct a solution to the equation using the theory of $\tau$ functions [17]. For this purpose, we first define the following real-valued functions for fixed $v$ and $y$.
(5.152) $\begin{split}&A_{n}\left(t\right)=\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right),\\ &B_{n}\left(t\right)=\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y};\frac{2i}{y}\right).\end{split}$
Then the potential function can be written as follows
(5.153) $\Theta_{n}\left(t\right)=4\arctan\frac{B_{n}}{A_{n}}=\frac{2}{i}\log\frac{A_{n}+iB_{n}}{A_{n}-iB_{n}}=\frac{2}{i}\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}},$
where $\widetilde{\tau}_{n}=A_{n}+iB_{n}$. $\widetilde{\tau}$ corresponds to the $\tau$ functions of the one-component KP hierarchy, while $A$ and $B$ correspond to the $\tau$ functions of the two-component KP hierarchy. Then we have the following lemma.
Lemma 5.18.
The functions $A$ and $B$ defined by (5.152) satisfy the following equations.
(5.154) $D_{t}A_{n+1}\cdot A_{n}+D_{t}B_{n+1}\cdot B_{n}=\frac{L_{1}}{2y}\left(\beta_{9}-1\right)\left(A_{n+1}A_{n}+B_{n+1}B_{n}\right),$
(5.155) $D_{t}A_{n+1}\cdot B_{n}-D_{t}B_{n+1}\cdot A_{n}=\frac{L_{1}}{2y}\left(\beta_{9}+1\right)\left(A_{n+1}B_{n}-B_{n+1}A_{n}\right),$
where
(5.156) $J_{1}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)},\quad J_{2}(v,y)=\frac{\vartheta^{{}^{\prime}}_{1}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}\left(0;\frac{2i}{y}\right)},$
(5.157) $L_{1}(v,y)=d_{1}\left(\frac{v}{y};\frac{2}{y}\right)-d_{4}\left(\frac{v}{y};\frac{2}{y}\right)+J_{1}(v,y)+J_{2}(v,y),$
(5.158) $\beta_{9}=\frac{d_{1}\left(\frac{v}{y};\frac{2}{y}\right)+d_{4}\left(\frac{v}{y};\frac{2}{y}\right)-J_{1}(v,y)+J_{2}(v,y)}{L_{1}(v,y)}.$
Proof.
[…]
by $h$ and putting $h=0$, we obtain
(5.164) $\begin{split}&D_{t}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{4}\left(\frac{v}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{4}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}J_{1}(v,y)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Also, after differentiating the equation in (A.7) with
[…]
by $h$ and putting $h=0$, we obtain
(5.170) $\begin{split}&D_{t}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{4}\left(\frac{v}{y}\right)}{\vartheta_{4}\left(\frac{v}{y}\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{4}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}J_{1}(v,y)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Also, after differentiating the equation in (A.7) with
[…]
by $h$ and putting $h=0$, we obtain
(5.176) $\begin{split}&D_{t}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{v}{y}\right)}{\vartheta_{1}\left(\frac{v}{y}\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{1}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}J_{2}(v,y)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Also, after differentiating the equation in (A.7) with
[…]
by $h$ and putting $h=0$, we obtain
(5.182) $\begin{split}&D_{t}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\cdot\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}\frac{\vartheta^{{}^{\prime}}_{1}\left(\frac{v}{y}\right)}{\vartheta_{1}\left(\frac{v}{y}\right)}\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}\frac{\vartheta_{4}\left(\frac{v}{y}\right)\vartheta^{{}^{\prime}}_{1}\left(0\right)}{\vartheta_{1}\left(\frac{v}{y}\right)\vartheta_{4}\left(0\right)}\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &=\frac{1}{y}d_{1}\left(\frac{v}{y};\frac{2}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{1}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right)\\ &-\frac{1}{y}J_{2}(v,y)\vartheta_{1}\left(\frac{v}{y}\left(n+\frac{1}{2}\right)+\frac{t}{y}\right)\vartheta_{4}\left(\frac{v}{y}\left(n-\frac{1}{2}\right)+\frac{t}{y}\right).\end{split}$
Adding (5.164) and (5.170) yields (5.154), and subtracting (5.176) from (5.182) yields (5.155). ∎
Then we have the following theorem.
@@
(5.183) $\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)$ $=\left(1+\cos\lambda\right)\rho_{+}\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right),$
(5.184) $\frac{d}{dt}\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)$ $=\left(1-\cos\lambda\right)\rho_{-}\sin\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right).$
Proof.
We first show that the potential function satisfies
@@
using Lemma 5.18. Then we show that
(5.186) $\frac{L_{1}(v,y)}{y}\frac{1}{1+\cos\lambda}=i\frac{\vartheta_{1}^{{}^{\prime}}(0)}{\vartheta_{1}(iv)}\frac{u_{1}}{\vartheta_{3}(r)^{2}\vartheta_{1}(r)^{2}}=\rho_{+}.$
Suppose that a solution to the equation (5.185) is expressed as (5.153). Then both sides of the equation can be rewritten as follows.
@@
(5.187) $\begin{split}&\frac{d}{dt}\left(\frac{\Theta_{n+1}+\Theta_{n}}{2}\right)=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n+1}}+\log\frac{\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n}}\right)\\ &=\frac{1}{i}\frac{d}{dt}\left(\log\frac{\widetilde{\tau}_{n+1}}{\widetilde{\tau}^{*}_{n}}-\log\frac{\widetilde{\tau}^{*}_{n+1}}{\widetilde{\tau}_{n}}\right)\\ &=\frac{1}{i}\left(\frac{D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}-\frac{D_{t}\widetilde{\tau}^{*}_{n+1}\cdot\widetilde{\tau}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}\right),\end{split}$
and
(5.188) $\begin{split}&\sin\left(\frac{\Theta_{n+1}-\Theta_{n}}{2}\right)=\frac{1}{2i}\left(\frac{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}-\frac{\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}}{\widetilde{\tau}_{n+1}\widetilde{\tau}^{*}_{n}}\right).\end{split}$
Then it can be seen that a sufficient condition for (5.153) to be a solution of (5.185) is that
(5.189) $\begin{split}&D_{t}\widetilde{\tau}_{n+1}\cdot\widetilde{\tau}^{*}_{n}=-\frac{L_{1}}{2y}\widetilde{\tau}^{*}_{n+1}\widetilde{\tau}_{n}+\frac{L_{1}}{2y}\beta_{9}\widetilde{\tau}_{n+1}\widetilde{\tau}_{n}^{*},\end{split}$
is satisfied. Substituting $\widetilde{\tau}_{n}=A_{n}+iB_{n}$ for this, we get
(5.190) $\begin{split}&D_{t}\left(A_{n+1}\cdot A_{n}+B_{n+1}\cdot B_{n}\right)-iD_{t}\left(A_{n+1}\cdot B_{n}-B_{n+1}\cdot A_{n}\right)\\ &=\frac{L_{1}}{2y}\left(\beta_{9}-1\right)\left(A_{n+1}A_{n}+B_{n+1}B_{n}\right)-i\frac{L_{1}}{2y}\left(\beta_{9}+1\right)\left(A_{n+1}B_{n}-B_{n+1}A_{n}\right).\end{split}$
The real part of this equation yields (5.154), and the imaginary part yields (5.155). Thus, Lemma 5.18 shows that the potential function satisfies the equation (5.185). We then calculate $L_{1}(v,y)/y$. From (A.15), (A.16) and (A.20), we see
(5.191) $\begin{split}\frac{L_{1}(v,y)}{y}&=\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)-\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}^{{}^{\prime}}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}\\ &+\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}+\frac{1}{y}\frac{\vartheta_{1}^{{}^{\prime}}\left(0;\frac{2i}{y}\right)\vartheta_{4}\left(\frac{v}{y};\frac{2i}{y}\right)}{\vartheta_{4}\left(0;\frac{2i}{y}\right)\vartheta_{1}\left(\frac{v}{y};\frac{2i}{y}\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}+i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{2}\left(0;iy\right)}\frac{\vartheta_{2}\left(iv;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\\ &=i\frac{\vartheta_{1}^{{}^{\prime}}\left(0;iy\right)}{\vartheta_{1}\left(iv;iy\right)}\frac{\vartheta_{4}\left(iv;iy\right)\vartheta_{2}\left(0;iy\right)+\vartheta_{2}\left(iv;iy\right)\vartheta_{4}\left(0;iy\right)}{\vartheta_{4}\left(0;iy\right)\vartheta_{2}\left(0;iy\right)}.\end{split}$
On the other hand, we see
(5.192) $\begin{split}&1+\cos\lambda\\ &=\frac{\left\{\vartheta_{1}\left(\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(-\frac{iv}{2}+r;iy\right)+\vartheta_{1}\left(-\frac{iv}{2}+r;iy\right)\vartheta_{3}\left(\frac{iv}{2}+r;iy\right)\right\}^{2}}{2u_{1}}.\end{split}$
We can see that (5.186) holds by rewriting (5.191) and (5.192) using the following.
(5.193) $\begin{split}&\vartheta_{2}\left(iv/2\right)\vartheta_{4}\left(iv/2\right)\left\{\vartheta_{1}\left(iv/2+r\right)\vartheta_{3}\left(-iv/2+r\right)+\vartheta_{1}\left(-iv/2+r\right)\vartheta_{3}\left(iv/2+r\right)\right\}\\ &=\vartheta_{3}\left(r\right)\vartheta_{1}\left(r\right)\left\{\vartheta_{4}\left(iv\right)\vartheta_{2}\left(0\right)+\vartheta_{2}\left(iv\right)\vartheta_{4}\left(0\right)\right\},\\ &\vartheta_{3}\left(r\right)\vartheta_{1}\left(r\right)\left\{\vartheta_{1}\left(iv/2+r\right)\vartheta_{3}\left(-iv/2+r\right)+\vartheta_{1}\left(-iv/2+r\right)\vartheta_{3}\left(iv/2+r\right)\right\}\\ &=\vartheta_{2}\left(iv/2\right)\vartheta_{4}\left(iv/2\right)\left\{\vartheta_{4}\left(2r\right)\vartheta_{2}\left(0\right)-\vartheta_{2}\left(2r\right)\vartheta_{4}\left(0\right)\right\},\\ &\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)\left\{\vartheta_{4}\left(2r\right)\vartheta_{2}\left(0\right)-\vartheta_{2}\left(2r\right)\vartheta_{4}\left(0\right)\right\}\\ &=2\vartheta_{3}\left(r\right)^{2}\vartheta_{1}\left(r\right)^{2}.\end{split}$
@@
(5.194) $\frac{d}{dt}\Theta_{n}=4\beta_{1}\left(\eta_{1}\cos\lambda+\eta_{2}\sin\lambda\right)\tan\left(\frac{\Theta_{n+1}-\Theta_{n-1}}{4}\right),$
@@
which is also called the semi-discrete potential mKdV equation [16, 15, 7]. In other words, the potential function $\Theta$ defined by (5.149) is a simultaneous solution of three differential-difference equations, (5.183), (5.184) and (5.194). It does not usually happen that these equations are simultaneously solved. For example, it can be checked that the 2- and 3-soliton solutions of the semi-discrete sine-Gordon equation constructed in [7, 8] are not solutions of the semi-discrete potential mKdV equation.